# Characterizing and Efficiently Accelerating Multimodal Generation Model Inference

### v1

@@ V Key Lessons and Concluding Remarks
Generative AI technologies are reshaping the computing landscape by offering new capabilities. This paper presents detailed system performance characterization of key multimodal generation models from Meta: Llama, Seamless, Chameleon, and gDLRM. We highlight the distinct resource requirements and performance characteristics, emphasizing the need for tailored optimization strategies. Enabling state-of-the-art system-level optimizations, such as Flash Attention/SDPA, torch.compile, CUDA Graph and quantization, brings significant improvements, leading to stronger baseline performance. Enabling workload-specific optimizations, such as LayerSkip, unlocks even further optimization opportunities by exploiting workload specific characteristics.
Our analysis reveals several key insights that are crucial for the computer architecture community:
@@
  • Multi-modal models exhibit unique workload characteristics that set them apart from traditional AI models. Our quantitative results demonstrate significant difference in latency, compute and memory requirements across different modalities and tasks. For instance, we observed that Text to Image and Image-Text to Text tasks of Chameleon demands 1.7$\times$ more compute than HSTU, while the arithmetic intensity of HSTU is 1.25$\times$ higher.
@@
  • Optimization solutions must consider the entire model inference pipeline end-to-end. Our research shows that focusing on isolated components may lead to suboptimal performance gains. For example, optimizing only the attention operation with SDPA gives 1.43$\times$ improvement, and additionally optimizing linear operations with AutoQuant gives additional 3.06$\times$ speedup, resulting in total 4.38$\times$ inference time speedup.
@@
  • While new hardware accelerators are exciting prospects, our results emphasize the importance of first exhausting state-of-the-art software optimizations. We demonstrated that enabling state-of-the-art optimization, such as, SDPA, torch.compile, and AutoQuant, already led to an 4.38$\times$ performance improvement across all the models, highlighting the untapped potential in existing hardware.
@@
  • The diversity in model architectures, input modalities necessitates flexible and adaptable optimization strategies. Our work shows that a one-size-fits-all approach is insufficient, as evidenced by showing that PyTorch SDPA might not be always helpful depending on the significance of the attention operation portion of the total runtime.
@@
VI Acknowledgement
This work is an outcome of the extensive collaborations with many teams: Chameleon, Seamless, and HSTU. We are thankful for the valuable insights, numerous discussions, and refinement on the multimodal models. We would also like to thank the PyTorch and the xFormers teams, especially their inputs on ML system optimization. In particular, we thank Mostafa Elhoushi for the assistance in implementing and debugging LayerSkip for the multi-modal generation tasks.

### v2

@@ 5 Key Lessons and Concluding Remarks
Generative AI technologies are reshaping the computing landscape by offering new capabilities. This paper characterizes system performance of key multimodal models from Meta: Llama, Seamless, Chameleon, and gDLRM and emphasizes distinct resource needs and performance patterns requiring tailored optimizations. Enabling state-of-the-art system-level optimizations, such as Flash Attention/SDPA, torch.compile, CUDA Graph and quantization, strengthen baseline performance. Enabling workload-specific optimizations unlocks even further optimization opportunities by exploiting workload specific characteristics. We present reveals key insights for the computer architecture community:
@@
  • Multi-modal models show distinct workload patterns compared to traditional AI models. Our quantitative results demonstrate difference in latency, compute and memory requirements across modalities and tasks. For instance, we observed that T-I and IT-T tasks of Chameleon demands 1.7$\times$ more compute than HSTU, while the arithmetic intensity of HSTU is 1.25$\times$ higher.
@@
  • Optimization solutions must consider the whole end-to-end inference pipeline. Our research shows that focusing on isolated components may lead to suboptimal performance gains. For example, optimizing only the attention operation with SDPA gives 1.43$\times$ improvement, and additionally optimizing linear operations with AutoQuant gives additional 3.06$\times$ speedup, resulting in total 4.38$\times$ inference speedup.
@@
  • While new hardware accelerators are exciting prospects, our results emphasize the importance of first exhausting state-of-the-art software optimizations. We demonstrated that enabling state-of-the-art optimization, SDPA, torch.compile, and AutoQuant, led to an 4.38$\times$ performance improvement across all the models, highlighting the untapped potential in existing hardware.
@@
  • The diversity in model architectures, modalities necessitates flexible and adaptable optimization strategies. PyTorch SDPA’s effectiveness varies with attention operation’s proportion of total runtime, showing one-size-fits-all approach is insufficient.
@@
6 Acknowledgment
This work is an outcome of the extensive collaborations with many teams: Chameleon, Seamless, and HSTU. We are thankful for the valuable insights, numerous discussions, and refinement on the multimodal models. We would also like to thank the PyTorch and the xFormers teams, especially their inputs on ML system optimization.